OpenStudy (anonymous):

A new cruise ship line launched 3 ships. The PP, CP, and MP. The CP has 31 more rooms than the PP. The MP has 35 fewer rooms than three times the number of rooms on the PP. Find the number of rooms for each ship if the total number of rooms for the three ships is 606. PP rooms? CP rooms? MP rooms?

OpenStudy (anonymous):

let the no of rooms for PP ship is x then setting the equation for three ships x+(x+31)+(3x-35)=606

OpenStudy (anonymous):

can you solve for x?

OpenStudy (anonymous):

@candrie

OpenStudy (anonymous):

ok solving the above equation for x 5x-4=606 5x=606+4 5x=610 x=610/5 x=122 122 rooms on PP 122+31 =153 rooms on CP 3(122)-35=331 rooms on MP